OpenStudy (anonymous):

Justin has 6 bags of dog food. His dogs eat 3/4 bag each day. How long will his dog food last? A. 4 1/2 days B. 6 3/4 days C. 7 days D. 8 days

OpenStudy (anonymous):

how many times does 3/4 go into 6

OpenStudy (anonymous):

and then you will get your answer

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thx's

OpenStudy (anonymous):

do you understand or need more help?

OpenStudy (anonymous):

i understand now thank you

OpenStudy (calculusxy):

\[\huge 6 \div \frac{3}{4} = \frac{6}{1} \times \frac{4}{3} = \color{green}{8}\]

OpenStudy (kendricklamar2014):

@alex6799 please dont give direct answers. All of these users are trying to help @Crazy_questions understand this question so he/she can solve it. Giving him/her the answer will not help in any way.

OpenStudy (anonymous):

im a he BTW

OpenStudy (kendricklamar2014):

Thanks for letting me know :)